OpenStudy (anonymous):

Transform the absolute value inequality into a compound inequality |6x+7| is less than or equal to 5. can someone plz help me?

OpenStudy (jdoe0001):

\(\bf |6x+7| \le 5\implies \begin{cases} +(6x+7) \le 5\\ \quad \\ \bf -(6x+7) \le 5 \end{cases}\implies -5\le 6x+7 \le 5\)

OpenStudy (jdoe0001):

each absolute value expression has 2 scenarios, thus you'd end up with 2 values for "x" one for the positive scenario, and one for the negative scenario

OpenStudy (anonymous):

what is a compound inequality?

OpenStudy (jdoe0001):

compound = composed of more than 1 element

OpenStudy (anonymous):

so my answer would just be 6x+7 is less than or equal to 5 and 6x+7 is greater than or equal to -5 ?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

but thats just rewriting the equation into two mini equations.

OpenStudy (jdoe0001):

yes it's

OpenStudy (jdoe0001):

it's all you're asked to do though

OpenStudy (jdoe0001):

I mean, you can solve for "x" if you want. you'd get 2 values 1 for each scenario, but you're not asked to do so this time

OpenStudy (anonymous):

so then why did you write -5<6x+7<5? instead of two mini equations?

OpenStudy (anonymous):

with the equal signs included, sorry i didnt know how to write them

OpenStudy (jdoe0001):

hmm I did write the 2 "mini equations" just in a bit different form \(|6x+7| \le 5\implies \begin{cases} \color{blue}{+(6x+7) \le 5}\\ \quad \\ \color{blue}{-(6x+7) \le 5} \end{cases}\implies -5\le 6x+7 \le 5\)

OpenStudy (anonymous):

but why are the positive and negative signs not in front of the 5?

OpenStudy (anonymous):

and are we supposed to flip the equation for -5?

OpenStudy (jdoe0001):

yeap, because you're multiplying for a negative number, so \(\bf |6x+7| \le 5\\\implies \begin{cases} +(6x+7) \le 5\implies 6x+7\le 5\\ \quad \\ \bf -(6x+7) \le 5\implies -1(-(6x+7)) \le -1(5)\implies 6x+7\ge -5 \end{cases}\)

OpenStudy (jdoe0001):

shoot got ..truncated... \(\bf |6x+7| \le 5\\\quad \\ \begin{cases} +(6x+7) \le 5\implies 6x+7\le 5\\ \quad \\ \bf -(6x+7) \le 5\implies -1(-(6x+7)) \le -1(5)\implies 6x+7\ge -5 \end{cases}\)

OpenStudy (anonymous):

so I can just write 6x+7 < or euqal to 5 OR 6x+7> or equal to -5 or do I have to write it like: {x| 6x+7<5 or 6x+7>_5}

OpenStudy (jdoe0001):

either way looks fine to me, is just a matter of notation

OpenStudy (anonymous):

ohhh ok thanks so much!

OpenStudy (jdoe0001):

yw